OpenStudy (anonymous):

how do you find this without using a calculator:

OpenStudy (anonymous):

\[(4-x)\log _{7}49=x+1\]

sam (.sam.):

\[(4-x)\log_{7}(7^{2})=x+1\] 2(4-x)=x+1 8-2x=x+1 7=2x x=7/2

sam (.sam.):

Mistake there 2(4-x)=x+1 8-2x=x+1 7=3x x=7/3

OpenStudy (anonymous):

ok thanks!